A particle starts from 8.0 m at t_0 = 0 and moves
Answer #1

Part B: Informatio of part A is necessary.

Part C:
From the graph the acceleration is:

a = \frac{\Delta v}{\Delta t}= 4/1 = 4 m/s^2

the position of the object is:

x =x_0+v_0t +\frac{1}{2}t^2

x = 8.0-4.0t +2.0t^2

for each time:

x_2 = 8 m

x_3 = 14 m

x_4 = 24 m